The problem: Find two numbers whose sum is 44 and whose difference is 22
I already know the answer - I had no problem figuring it out. What I need to see is the proper way to solve it. I got the answers 33 and 11. This is how I started it:
x + y = 44
x - y = 22
x = 22 + y and x = 44 - y .....what next? I plugged numbers in until I found the answer but there must be an easier way?

Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

x + y = 44
x - y = 22
------------
Add the two equations to get:
2x = 66
x = 33
----------
Substitute into x+y=44 to solve for y:
33+y = 44
y = 11
==============
Cheers,
Stan H.

Answer by scott8148(6628)

if you add your two qriginal equations together, you get 2x=66 (the y's cancel each other)

so x=33; and 33+y=44 so y=11